OpenStudy (anonymous):

A rectangular coil with sides 0.140 m by 0.220 m has 493 turns of wire. It is rotated about its long axis in a uniform magnetic field of 0.560 T that is perpendicular to the long axis. At what frequency, in complete cycles per second, must the coil be rotated for it to generate a maximum voltage of 115 V?

OpenStudy (anonymous):

The best I've come up with is \[Emf =N \omega AB \sin \omega t\]\[\omega = f2\pi \]Which would give \[Emf =N f2\pi AB \sin f2\pi t\] But this still leaves me with two variables which makes me feel like I'm not using the correct equations. I'll keep searching, but if someone could let me know if I'm on the right track that would be awesome!

OpenStudy (irishboy123):

you need only the max E so you can ignore the sin function and go with Emax = NBAw where N = turns B = field A = area w = ang vel and w = 2πf (f = freq) f = (1/2π) Emax/NBA

OpenStudy (irishboy123):

soz - the funny stiff is because i tried to use the symbol for pi rathern than pi on my keyboard. clearly i mean w = 2 pi f etc etc

OpenStudy (anonymous):

The max EMF if what got me. Thanks for the help